# Physics meets Topology: Physics-informed topological neural networks for learning rigid body dynamics

### v1

@@ 3.1 Forward dynamics in increasingly complex scenarios
We conduct our evaluations using three benchmark video datasets of increasing complexity: MoVi-spheres, MoVi-A, and MoVi-B kubric . Each experiment consists of a video where a set of rigid objects interact within an environment with unknown characteristics – such as gravity and air resistance – that can only be inferred from observations. The MoVi-spheres dataset consists solely of spherical objects, while the MoVi-A includes a mixture of cubes, spheres, and cylinders. The most complex, MoVi-B, contains real-world everyday objects, such as cones and teapots. In each experiment, between 3 to 10 objects are randomly selected, each assigned varying physical parameters (e.g. size, mass, friction). The objects are initialized with random positions and velocities and then released onto a flat surface to simulate dynamic interactions. Detailed descriptions of the datasets are provided in Supplementary Information Sec. 2.1.
@@
We test our framework in an autoregressive manner: the predicted acceleration $\hat{\ddot{\bm{x}}}^{t}$ at time $t$ is used to compute the next position $\hat{\bm{x}}^{t+1}$, which is subsequently fed back as input to predict $\hat{\bm{x}}^{t+2}$, as elaborated in Sec. 2.1. To evaluate the framework’s performance over both short-term and long-term predictions, we consider different rollout horizons $T$ ranging from $25$ to $100$ steps. Our primary results are presented in Fig. 4 for a rollout horizon of $T=50$, aligning with the reported results by the current state-of-the-art baseline models FIGNet fignet and MeshGraphNet meshgraphnet , due to the absence of reproducible code for these models. For a more comprehensive analysis across various horizons $T=\{25,50,75,100\}$, please refer to Supplementary Information Sec. 3.1.
The results presented in Fig. 4 demonstrate that our framework effectively learns and predicts the dynamics of interacting rigid bodies with remarkable accuracy, closely aligning with the ground truth. In line with previous research studies, we report the average Root Mean Square Error (RMSE) for both object positions (Fig. 4a) and orientations (Fig. 4b) after 50 timesteps. On the MoVi-A dataset, our framework achieves high fidelity in capturing both the translational and rotational motions of rigid bodies. Moreover, on the more complex MoVi-B dataset, which includes a diverse set of real-world objects, our framework effectively handles collisions, maintaining robust performance despite the increased complexity. Across all datasets, our method consistently outperforms both state-of-the-art methods MeshGraphNet meshgraphnet and FIGNet fignet by a substantial margin, demonstrating the effectiveness of our higher-order representations and physics-informed inductive biases.
Importantly, our method is capable of performing $50\%$ more rollout steps before reaching the same translational and rotational errors. The best baseline reaches an RMSE of $0.115$ [m] at 50 steps on the MoVi-A dataset, whereas our framework reaches this value after 75 steps. This significant improvement allows more accurate predictions further into the future, supporting better and more robust decision-making in dynamic scenarios. For further analysis of extended rollout horizons, please refer to Sec. 3.1. in the Supplementary Information.
Figure 6: Counterfactual rollout example. (a) Original experiment444 The sequence shown in Fig. 4 is taken from experiment n°21 of the MoVi-A dataset. ; (b) Altered experiment with one sphere removed, shown as a black skeleton to ease understanding. With our dynamic and flexible higher-order topological representation of the environment, our model can easily accommodate any changes in initial conditions: position, orientation, velocity, material properties, …
@@
To further validate the robustness and precision of our framework, we present a challenging multi-object rollout example in Fig. 3. In this scenario, the final positions $\hat{\bm{x}}$ and orientations $\hat{\bm{q}}$ of all objects closely match the ground truth trajectories, highlighting the framework’s precision and robustness in handling complex interactions among multiple entities.
3.2 Counterfactual capabilities and generalization
In addition to the high accuracy of our framework, two critical benefits arise from our design choices. First, the higher-order topological representation of the environment facilitates the modification of initial conditions, enabling the exploration of counterfactual scenarios. Counterfactual scenarios refer to hypothetical situations that differ from the actual conditions, allowing researchers to assess potential outcomes by altering specific variables such as an object’s starting position or mass. These scenarios are inherently challenging because they require the model to accurately predict the effects of changes in initial conditions on the system’s subsequent dynamics, often involving complex and nonlinear interactions. For example, we can investigate questions like: "What if an object had a different starting position?" or "What if it had a different mass?". All scene parameters – including any object’s position, orientation, velocity, mass, and material properties – are fully configurable. Fig. 4 illustrates this flexibility with a counterfactual scenario where a large sphere, initially interacting with multiple objects, is removed at time $t_{0}$. Consequently, the dynamic interactions among the remaining objects evolve differently, leading to a distinct outcome. This counterfactual capability is particularly valuable for decision-making, as it enables the exploration of multiple hypotheses before committing to a specific course of action.
Building on this flexibility, our framework also demonstrates exceptional robustness and generalization capabilities. This second benefit is closely related to the robustness of our model and its capacity to handle out-of-distribution scenarios. By embedding physics-informed message passing into our framework, we enable it to learn fundamental physical laws and effectively extrapolate to unseen complex object geometries. To validate this capability, we trained three models – S, A, and B – on the MoVi-spheres, MoVi-A, and MoVi-B datasets, respectively. Each model was exposed to a different level of object complexity: S only to spheres, A to basic shapes (including spheres), and B to real-world objects (including basing shapes). We then evaluated each model’s generalization performance across all datasets. The results, summarized in Fig. 7, demonstrate that HOPNet generalizes very effectively to unseen object geometries. Despite being trained solely on spherical objects, model S achieves competitive and even superior performance on MoVi-A and MoVi-B than models A and B (Fig. 7a and 7b). This exceptional generalization ability can be attributed to model’s S exposure to a more diverse range of collision scenarios. In contrast to models A and B, which primarily encountered static and slow-speed interactions, model S was exposed to more dynamic and varied collisions (Fig. 7c). This difference in variety can be explained by the nature of each dataset’s objects: spheres remain in continuous motion, while more complex shapes like cubes or cylinders tend to stabilize on one of their faces, leading to fewer and more similar interactions overall.
In summary, this analysis demonstrates HOPNet’s superior ability to generalize to unseen complex geometries and emphasizes the importance of having a large variety of operating conditions in a dataset. An example of this generalization is presented in Supplementary Information Sec. 3.2.
Figure 7: Generalization performance of HOPNet. (a) Translation and (b) orientation errors after $T=50$ rollout steps of three different models – S, A, and B – trained on MoVi-spheres, MoVi-A, and MoVi-B respectively. Error bars indicate the minimum, mean, and maximum performance across three independent random seeds. (c) Distribution of the collision velocities encountered by each model during training. Model S, trained solely on spherical objects, generalizes effectively to more complex object geometries in MoVi-A and MoVi-B. This ability arises from model S’s exposure to a wider range of collision dynamics compared to the slower and more static collisions seen by models A and B. Further details about the datasets are available in Supplementary Information Sec. 2 and 3.
@@ 3.3 Importance of physical consistency
To further demonstrate the impact of our core contributions, we conducted ablation studies on both the higher-order topological representation and the physics-informed inductive biases. For the topological representation, we compared our full higher-order representation against a version that omits the higher-order object-level cells $X^{(4)}$. We did not explicitly test the removal of other higher-order cells, as existing baselines MeshGraphNet meshgraphnet and FIGNet fignet can be considered approximations of reduced representations without mesh triangles $X^{(2)}$ and object-level cells $X^{(4)}$, respectively. To assess the importance of physics-informed biases, we performed two key modifications:(1) omitting key information inside the node features $\bm{h}^{(0)}$, specifically the relative distance between the node and the object’s center of mass, and (2) removing the multi-step sequential message-passing. Further architectural details about the models used in these ablation experiments are provided in Supplementary Information Sec. 3.3.
When comparing our full combinatorial complex representation to its ablated versions, Fig. 8a reveals that removing the object-level cells $X^{(4)}$ leads to a significant drop in performance, with a translation RMSE $44.4\%$ larger than the default HOPNet framework. This decline is expected since intra-object messages are restricted to node-to-node communication, which diminishes the model’s expressiveness and fails to capture the rigidity of each object. Although a virtual node is added at the center of each object to enhance multi-hop communication between nodes, the lack of physical consistency results in a noticeable reduction in accuracy. This modified version resembles FIGNet fignet in its features but still retains the physics-informed message-passing from HOPNet. With the performance of this ablated model ($\text{RMSE}^{\text{pos}}=0.078$ [m]) falling in between HOPNet ($\text{RMSE}^{\text{pos}}=0.054$ [m]) and FIGNet ($\text{RMSE}^{\text{pos}}=0.115$ [m]), this study also demonstrate the benefits of HOPNet’s physics-informed message-passing compared to a standard convolutional one.
Beyond our topological representation, our second key contribution is the incorporation of physics-informed biases within our topological neural network. To underscore the importance of feature selection, we tested the removal of a critical attribute from the node cell features $\bm{h}^{(0)}$: the relative distance between the node and the object’s center of mass. This feature is particularly important from a physics standpoint, as it provides structural information about the object and enables the prediction of how collisions near specific nodes affect the object’s overall momentum. To evaluate the significance of this feature, we tested our full framework and its simplified version on two datasets: MoVi-spheres and MoVi-A. In the MoVi-spheres dataset, which consists exclusively of spherical objects, removing the "center-mass distance" feature is expected to have a minimal impact, as a sphere’s moment of inertia is uniform in all directions. In contrast, the cubes and cylinders in MoVi-A possess anisotropic moments of inertia, making the nodes’ distances to the center of mass critical for accurately predicting collisions with these objects.
The results presented in Fig. 8b validate our hypothesis. We only observe a slight performance decline on the MoVi-spheres dataset with $\text{RMSE}^{\text{pos}}$ increasing from $0.034$ [m] to $0.040$ [m], primarily due to the coarse resolution of the spherical meshes. In contrast, there is a substantial drop in accuracy on the MoVi-A dataset, highlighting the critical importance of physics-informed biases for accurately modeling objects with more complex shapes. The $\text{RMSE}^{\text{pos}}$ becomes $26\%$ larger at $0.068$ [m] compared to $0.054$ [m] for the default model. Additionally, on both datasets, the seed-to-seed variability (shown by the error bars) greatly increases compared to the default HOPNet model, underscoring the robustness of our approach.
@@
Figure 8: Importance of physical consistency in HOPNet. (a) Ablation study on higher-order topological representations; (b) Ablation study on physics-based features; (c) Ablation study on physics-informed message-passing. Error bars indicate the minimum, mean, and maximum performance across three independent random seeds.
Finally, we assess the impact of embedding physics-informed inductive biases within our message-passing scheme. Specifically, we compare our physics-informed sequential message-passing (illustrated in Fig. 2) with a non-sequential version, where information flows freely between all cells without any structural guidance. This alternative approach removes the inductive bias that constrains the transfer pathways within our topological neural network. The results, presented in Fig. 8c, clearly demonstrate that without these guided pathways, the model struggles to efficiently learn collision dynamics and has difficulty capturing the interactions among higher-order groups. The $\text{RMSE}^{\text{pos}}$ increases to $0.061$ [m], even though this variant of the model contains $34\%$ more learnable parameters than the default HOPNet. This highlights the critical role of structured information flow in modeling and learning complex physical interactions.
@@ 3.4 Hyperparameter robustness and efficiency
Our framework demonstrates significant robustness to variations in hyperparameter selection, highlighting its efficiency and generalizability across diverse applications. When testing a smaller variant of HOPNet by reducing the total number of learnable parameters by $74\%$ through a decreased hidden embedding size, we observed only a minor decrease in performance. This finding confirms the framework’s efficiency in learning rigid-body interactions. Additionally, increasing the total number of message-passing steps leads to improved accuracy, a trend consistent with conventional GNNs. A detailed analysis of model size effects is provided in Supplementary Information Sec. 3.4.
Moreover, our framework shows minimal sensitivity to critical modeling hyperparameters, such as the collision radius $d_{c}$ used to define collision contact cells $X^{(3)}$. Performance remained stable even when the collision radius was halved or doubled, further validating the method’s stability across a wide range of hyperparameter values. This robustness minimizes the need for extensive fine-tuning, thereby broadening our framework’s applicability across different tasks and scenarios. Additional details on this analysis can be found in Supplementary Information Sec. 3.5.
@@ 4 Discussion
We validated our approach on a range of increasingly complex datasets, from simple shapes to real-world objects with intricate geometries. In addition to outperforming baseline models, our framework demonstrated robust generalization to unseen scenarios and the ability to address counterfactual questions – an essential feature for decision-making in dynamic systems. Although this work focuses on rigid-body dynamics, the framework is adaptable enough to extend to deformable objects, such as soft bodies or fluids, and to incorporate additional forces like magnetic or electric fields.
@@
Nevertheless, scaling our method to extremely complex meshes with millions of faces remains a challenge. While downsampling or mesh approximation could mitigate the computational load, alternative methods using implicit surface modeling might be better suited for such extreme cases.
Finally, while our framework currently relies on explicit state information (e.g. geometry, mass, friction) for counterfactual reasoning and generalization, possible extensions could infer these parameters implicitly from data – past observations – in other environments. Additionally, coupling our model with visual perception systems is a promising direction for further exploration, enabling learning directly from raw pixel data.

### v3

@@ 3.1 Forward dynamics in increasingly complex scenarios
Figure 6: Generalization performance of HOPNet. (a) Translation and (b) orientation errors after $T=50$ rollout steps of three different models (S, A, and B) trained on MOVi-spheres, MOVi-A, and MOVi-B respectively. Error bars indicate the mean and standard deviation across three independent random seeds. (c) Distribution of the collision velocities encountered by each model during training. Model S, trained solely on spherical objects, generalizes effectively to more complex object geometries in MOVi-A and MOVi-B. This ability arises from model S’s exposure to a wider range of collision dynamics compared to the slower and more static collisions seen by models A and B. Further details about the datasets are available in Supplementary Information Sections 2 and 3.
@@
We conduct our evaluations using four benchmark video datasets of increasing complexity: MOVi-spheres, MOVi-A, MOVi-B and MOVi-C kubric . Each experiment consists of a video where a set of rigid objects interact within an environment with unknown characteristics—such as gravity and air resistance—that can only be inferred from observations. The MOVi-spheres dataset consists solely of spherical objects, while the MOVi-A includes a mixture of cubes, spheres, and cylinders with coarse mesh resolutions (51 to 64 nodes per object). MOVi-B increases complexity by incorporating real-world everyday objects, such as cones and teapots, with meshes containing up to 1590 nodes. The most challenging dataset, MOVi-C, uses high-resolution 3D-scanned household items from Google’s Scanned Objects dataset downs2022google averaging 7917 nodes per object and reaching up to 50561 nodes.
Each experiment contains between 3 and 10 randomly selected objects with varying physical parameters (i.e., mass, friction) and sizes. In MOVi-A and MOVi-B, objects are initially defined with an average dimension of 1 [m] and then assigned a scaling factor of either 0.7 or 1.4. In contrast, objects in MOVi-C retain their real-world sizes but are further scaled by a randomly sampled factor in the range [0.75,3.0][0.75,3.0][ 0.75 , 3.0 ]. Objects are initialized with random positions and velocities and then released onto a flat surface to undergo dynamic interactions. Further details about the datasets are provided in Supplementary Information Section 2.1.
@@
We test our framework in an autoregressive manner: the predicted acceleration $\hat{\ddot{\mathbf{x}}}^{t}$ at time $t$ is used to compute the next position $\hat{\mathbf{x}}^{t+1}$, which is subsequently fed back as input to predict $\hat{\mathbf{x}}^{t+2}$, as elaborated in Section 2.1. To evaluate the framework’s performance over both short-term and long-term predictions, we consider different rollout horizons $T$ ranging from 252525 to 100100100 steps. Our primary results are presented in Figure 4 for a rollout horizon of $T=50$, aligning with the reported results by the current state-of-the-art baseline models FIGNet fignet and MeshGraphNet meshgraphnet . For a more comprehensive analysis across various horizons $T=\{25,50,75,100\}$, please refer to Supplementary Information Section 3.1.
Due to the absence of official reproducible code, we report results for two variants of FIGNet: the numbers published in the original paper fignet , and a ${}^{\text{reimpl}}$ trained on exactly the same datasets than HOPNet. The reimplementation is based on publicly available code, which we extended to support the MOVi datasets (see Supplementary Information Section 2.5 for details).
[…]
@@
To further validate the robustness and precision of our framework, we present a challenging multi-object rollout example in Figure 1. In this scenario, the final positions $\{\hat{\mathbf{x}}_{k}\}_{i=1...K}$ and quaternionic orientations $\{\hat{\mathbf{q}}_{k}\}_{i=1...K}$ of all $K$ objects closely match the ground truth trajectories, highlighting the framework’s precision and robustness in handling complex interactions among multiple entities.
3.2 Out-of-distribution generalization to unseen complex geometries
Figure 7: Extreme OOD generalization to complex real-world meshes. (a) Translation and (b) orientation errors after $T=50$ rollout steps on the MOVi-C dataset of three different models (S, A, and B) trained on MOVi-spheres, MOVi-A, and MOVi-B respectively. Error bars indicate the mean and standard deviation across three independent random seeds.
Beyond achieving high accuracy, HOPNet demonstrates strong Out-Of-Distribution (OOD) generalization to object geometries not encountered during training. This capability arises from the synergy between its physics-informed architecture and higher-order topological representation, which together allow the model to learn generalizable face-face collision dynamics rather than overfitting to object-specific patterns.
We distinguish between two forms of generalization: in-distribution and out-of-distribution (OOD). In-distribution generalization refers to variations within the training distribution, such as different object instances, object counts, or velocity profiles that are already represented in the training data. This type of generalization is typically assessed through standard test-set evaluation and is generally handled well by existing graph-based simulators battaglia2016interaction . In contrast, OOD generalization involves more substantial deviations from the training distribution—such as entirely novel object geometries, significantly higher object counts, or previously unseen velocity ranges. Successfully handling these scenarios requires models to capture the underlying physical principles, rather than relying on statistical patterns learned from the training data.
OOD generalization to novel geometries  To evaluate HOPNet’s ability to generalize to unseen object geometries, we trained three models (S, A, and B) on the MOVi-spheres, MOVi-A, and MOVi-B datasets, respectively. These datasets span increasing object complexity: S was trained only on spheres, A on basic shapes (including spheres), and B on real-world objects (including basic shapes). All three models were evaluated across all test sets. The results, shown in Figure 6, reveal that all three models generalize well, with relatively small differences in performance, even when models are evaluated far outside their training domain.
Notably, model S, trained exclusively on spherical objects, achieves comparable or superior performance on MOVi-A and MOVi-B compared to models A and B (Fig. 7a and 7b). This counterintuitive results can be explained by the diversity of collisions in the MOVi-spheres dataset. Because spheres roll and collide continuously, model S is exposed to a rich set of collisions with varying contact angles and speeds. In contrast, objects in MOVi-A and MOVi-B often stabilize on flat surfaces or sharp corners, leading to more redundant interactions (Fig. 7c). To validate this hypothesis, we modified model B’s training by masking slow collisions, reducing the bias toward near-static collisions. The resulting model outperforms the original models B and S, confirming that mitigating the bias toward static contacts leads to better generalization. These findings further support our hypothesis that exposure to diverse, dynamic collisions is critical for learning transferable collision dynamics. Details are provided in Supplementary Information Section 3.3.
Extreme OOD generalization to real-world meshes  To further assess HOPNet’s OOD generalization, we evaluated all three models on MOVi-C, a dataset featuring high-resolution 3D-scanned objects downs2022google . Compared to their respective training sets, MOVi-C scenes contain on average two orders of magnitude more mesh nodes, making this a highly challenging OOD generalization task.
Despite the substantial increase in geometric complexity, all models continue to perform remarkably well (Fig. 7). Consistent with the previous generalization task, model S achieves the lowest errors (0.1040.1040.104 [m] translation, 6.5646.5646.564 [deg] orientation), followed closely by model B (0.1050.1050.105 [m], 7.0947.0947.094 [deg]) and model A (0.1090.1090.109 [m], 6.9586.9586.958 [deg]).
The consistently strong performance across all HOPNet models (S, A, and B) reinforces our hypothesis that learning face-to-face collisions with diverse angles and velocities is sufficient for HOPNet to capture universal collision dynamics, independent of specific geometries or mesh resolutions.
[…]
Figure 8: Counterfactual rollout example. (a) Original experiment22footnotemark: 2 ; (b) Altered experiment with one sphere removed, shown as a black skeleton to ease understanding. With our dynamic and flexible higher-order topological representation of the environment, our model can easily accommodate any changes in initial conditions: position, orientation, velocity, material properties, … The illustrated sequence is taken from experiment 21 of the MOVi-A dataset.
3.3 Counterfactual capabilities
HOPNet’s accuracy and OOD generalization enable robust counterfactual reasoning, which is essential for planning, analysis, and decision-making. Counterfactual scenarios involve modifying specific scene parameters—such as object positions, velocities, or the presence of certain elements—to simulate hypothetical outcomes. While many graph-based simulators support such interventions in principle, achieving accurate long-horizon rollouts under novel conditions remains a significant challenge.
[…]
Figure 2 illustrates a concrete counterfactual example: a large sphere is removed from the scene at time $t_{0}$, and HOPNet accurately predicts the resulting evolution of the modified scene. This ability to simulate targeted interventions across diverse object configurations and physical conditions demonstrate the flexibility and practical utility of our approach.
@@ 3.4 Importance of physical consistency
Figure 9: Importance of physical consistency in HOPNet. (a) Ablation study on higher-order topological representations; (b) Ablation study on physics-based features; (c) Ablation study on physics-informed message-passing. Error bars indicate the mean and standard deviation across three independent random seeds.
To further demonstrate the impact of our core contributions, we conducted ablation studies on both the higher-order topological representation and the physics-informed inductive biases. For the topological representation, we compared our full higher-order representation against a version that omits the higher-order object-level cells $X^{(4)}$. We did not explicitly test the removal of other higher-order cells, as existing baselines MeshGraphNet meshgraphnet and FIGNet fignet can be considered as approximations of reduced representations without mesh triangles $X^{(2)}$ and object-level cells $X^{(4)}$, respectively. To assess the importance of physics-informed biases, we performed two key modifications:(1) omitting key information inside the node features $\mathbf{h}^{(0)}$, specifically the relative distance between the node and the object’s center of mass, and (2) removing the multi-step sequential message-passing. Further architectural details about the models used in these ablation experiments are provided in Supplementary Information Section 3.4.
When comparing our full combinatorial complex representation to its ablated versions, Figure 8a reveals that removing the object-level cells $X^{(4)}$ leads to a significant drop in performance, with a translation RMSE $4\%4$ larger than the default HOPNet framework. This decline is expected since intra-object messages are restricted to node-to-node communication, which diminishes the model’s expressiveness and fails to capture the rigidity of each object. Although a virtual node is added at the center of each object to enhance multi-hop communication between nodes, the lack of physical consistency results in a noticeable reduction in accuracy. This modified version resembles FIGNet fignet in its features but still retains the physics-informed message-passing from HOPNet. With the performance of this ablated model ($\text{RMSE}^{\text{pos}}=0.078$ [m]) falling in between HOPNet ($\text{RMSE}^{\text{pos}}=0.054$ [m]) and FIGNet ($\text{RMSE}^{\text{pos}}=0.115$ [m]), this study also demonstrates the benefits of HOPNet’s physics-informed message-passing compared to a standard convolutional one.
Beyond our topological representation, our second key contribution is the incorporation of physics-informed biases within our topological neural network. To underscore the importance of feature selection, we tested the removal of a critical attribute from the node cell features $\mathbf{h}^{(0)}$: the relative distance between the node and the object’s center of mass. This feature is particularly important from a physics standpoint, as it provides structural information about the object and enables the prediction of how collisions near specific nodes affect the object’s overall momentum. To evaluate the significance of this feature, we tested our full framework and its simplified version on two datasets: MOVi-spheres and MOVi-A. In the MOVi-spheres dataset, which consists exclusively of spherical objects, removing the "center-mass distance" feature is expected to have a minimal impact, as a sphere’s moment of inertia is uniform in all directions. In contrast, the cubes and cylinders in MOVi-A possess anisotropic moments of inertia, making the nodes’ distances to the center of mass critical for accurately predicting collisions with these objects.
[…]
@@ 3.5 Hyperparameter robustness and efficiency
[…]
Moreover, our framework shows minimal sensitivity to critical modeling hyperparameters, such as the collision radius $d_{c}$ used to define collision contact cells $X^{(3)}$. Performance remained stable even when the collision radius was halved or doubled, further validating the method’s stability across a wide range of hyperparameter values. This robustness minimizes the need for extensive fine-tuning, thereby broadening our framework’s applicability across different tasks and scenarios. Additional details on this analysis can be found in Supplementary Information Section 3.6.
@@ 4 Discussion
We validated our approach on a range of increasingly complex datasets, from simple shapes to real-world objects with intricate geometries. In addition to outperforming baseline models, our framework demonstrated robust generalization to unseen scenarios and the ability to address counterfactual questions—an essential feature for decision-making in dynamic systems. Although this work focuses on rigid-body dynamics, the framework is adaptable enough to extend to deformable objects, such as soft bodies or fluids, and to incorporate additional forces like magnetic or electric fields.
@@
Scaling our method to extremely complex meshes with millions of faces remains a challenge, primarily due to the computational cost of message passing over large topological structures. Mesh simplification techniques, such as quadric error decimationgarland1997surface , can reduce the complexity but often at the cost of losing geometric details. Adaptive mesh refinement narain2012adaptive offers greater flexibility by dynamically increasing resolution in regions of interest, for example around collision events. Hierarchical message passing, which has recently shown promise in fluid simulations bleeker2025neuralcfd , could further improve efficiency by enabling multi-scale information propagation. While implicit surface models avoid explicit meshing altogether, their applicability to dynamic, physics-driven scenes remains an open question. In future work, we aim explore these directions to improve scalability without compromising physical fidelity.
Finally, while our framework currently relies on explicit state information (e.g., geometry, mass, friction) for counterfactual reasoning and generalization, possible extensions could infer these parameters implicitly from data (past observations) in other environments. Additionally, coupling our model with visual perception systems is a promising direction for further exploration, enabling learning directly from raw pixel data.